rewrite the equation so that Y is a function of X:
4x-3(y-2)=15+y

Answer by checkley77(12844)

4x-3(y-2)=15+y
4x-3y+6=15+y
4x=y+3y=15-6
4x=4y+9
x=4y/4+9/4
x=y+2.25 ans.